# SEDM: Scalable Self-Evolving Distributed Memory for Agents

### v1

@@ Abstract
Long-term multi-agent systems inevitably generate vast amounts of trajectories and historical interactions, which makes efficient memory management essential for both performance and scalability. Existing methods typically depend on vector retrieval and hierarchical storage, yet they are prone to noise accumulation, uncontrolled memory expansion, and limited generalization across domains. To address these challenges, we present SEDM, Self-Evolving Distributed Memory, a verifiable and adaptive framework that transforms memory from a passive repository into an active, self-optimizing component. SEDM integrates verifiable write admission based on reproducible replay, a self-scheduling memory controller that dynamically ranks and consolidates entries according to empirical utility, and cross-domain knowledge diffusion that abstracts reusable insights to support transfer across heterogeneous tasks. Evaluations on benchmark datasets demonstrate that SEDM improves reasoning accuracy while reducing token overhead compared with strong memory baselines, and further enables knowledge distilled from fact verification to enhance multi-hop reasoning. The results highlight SEDM as a scalable and sustainable memory mechanism for open-ended multi-agent collaboration. The code will be released in the later stage of this project.
@@ 1 Introduction
In recent years, the rapid development of large-scale multi-agent systems (MAS) Wooldridge (2009); Lowe et al. (2017); Foerster et al. (2016); Yang et al. (2018a); Buşoniu et al. (2008). has expanded their application in diverse domains, including collaborative reasoning, decision-making, and autonomous planning Shoham et al. (2007). A central challenge in these open-ended and long-term tasks is the ability of agents to manage, interpret, and reuse information accumulated from continuous interaction with both peers and the environment Yang et al. (2018a). Without effective memory mechanisms, the sheer scale of historical interactions can easily overwhelm computational resources and compromise decision quality Foerster et al. (2018).
@@
In open-ended and long-term multi-agent tasks, each agent relies on its past memories, the observed states of other agents, and the current environment to determine subsequent actions or responses Das et al. (2017). During continuous interaction between agents and their environment, the MAS gradually accumulates extensive logs of interactions, invocation trajectories, and high-level policy memories OpenAI (2023). Such overwhelming amounts of information directly impact the efficiency and cost of decision-making, often leading to higher monetary costs and longer contextual requirements for inference Shuster et al. (2020). Therefore, designing an efficient and sustainable memory mechanism has become a critical issue for modern long-term multi-agent systems.
@@
Current methods primarily adopt vector retrieval and hierarchical memory structures to manage storage and retrieval efficiently Johnson et al. (2019b). Vector retrieval Johnson et al. (2019a); Karpukhin et al. (2020); Izacard and Grave (2021); Lewis et al. (2020); Guo et al. (2016) leverages semantic similarity to identify relevant entries, while hierarchical organization arranges information in layered structures according to abstraction levels Rae et al. (2021). These approaches have shown promise in improving retrieval accuracy and managing memory scalability Chen et al. (2016). However, in complex collaborative multi-agent tasks, their effectiveness diminishes, as the underlying assumptions of stability and linear growth do not hold Wang et al. (2020). This gap between theoretical promise and practical performance highlights several critical limitations that hinder their long-term applicability.
@@
One major challenge is the inevitable accumulation of noise, which severely degrades retrieval quality Fan et al. (2018). As the memory size expands without constraint, the system faces exponentially increasing computational costs in both retrieval and context construction Izacard and Grave (2021). This not only reduces overall efficiency but also amplifies the interference caused by redundant information Lewis et al. (2020). In particular, the presence of low-value or semantically irrelevant entries dilutes the contribution of high-quality information in retrieval results, impairing downstream task performance and leading to measurable declines in metrics Chen et al. (2017). In addition, the cumulative noise effect increases response latency and accelerates the nonlinear consumption of computational and storage resources Karpukhin et al. (2020), ultimately threatening both scalability and stability in long-term MAS operations Press et al. (2020).
@@
To overcome these limitations, we introduce Scalable Self-Evolving Distributed Memory (SEDM), a framework that transforms memory from a passive repository into an adaptive, self-optimizing, and verifiable component for multi-agent systems. Unlike conventional designs that treat memory as a static store, SEDM continually refines knowledge to enhance learning and decision-making efficiency in dynamic task environments. It operationalizes memory as an active mechanism by integrating verifiability and continuous self-improvement into the memory lifecycle. At its core, memory items undergo a rigorous admission process based on self-contained execution contexts (SCECs), such as Docker and ReproZip Merkel (2014); Chirigati et al. (2016), which package all necessary information for environment-free replay and offline validation. This mechanism provides empirical evidence for utility at write time, ensuring that only useful, high-quality experiences enter the memory repository. Once admitted, memory items are dynamically managed by a self-scheduling controller and enhanced through cross-domain knowledge diffusion. The controller leverages admission-derived weights, combined with semantic similarity, to schedule retrieval-time usage without costly reranking, while consolidation and progressive evolution continuously refine the repository by promoting stable items, merging redundancies, and pruning harmful ones. Beyond single-task settings, SEDM abstracts reusable insights into general forms, enabling knowledge distilled in one domain to be safely transferred and re-validated in others. Together, these components establish a scalable and auditable memory mechanism that enhances reasoning accuracy, reduces overhead, and supports sustainable long-term multi-agent collaboration.
@@
We evaluate SEDM on two representative benchmarks, FEVER Thorne et al. (2018) for fact verification and HotpotQA Yang et al. (2018b) for multi-hop reasoning, comparing against no-memory and G-Memory baselines Zhang et al. (2025). The results show that SEDM consistently improves task accuracy while significantly reducing token overhead, thereby achieving a better balance between performance and efficiency. Ablation studies confirm that both the verifiable admission mechanism and the self-scheduling controller contribute progressively to this gain, with the latter playing a key role in constraining prompt growth without sacrificing accuracy. Furthermore, cross-domain evaluation demonstrates that memory distilled from one dataset can transfer to another, with factual knowledge from FEVER notably boosting performance on HotpotQA. These findings highlight SEDM as a scalable, adaptive, and generalizable memory framework for long-term multi-agent reasoning.
@@
  • We propose Scalable Self-Evolving Distributed Memory (SEDM), a novel framework that transforms memory from a passive repository into an adaptive, verifiable, and continuously improving component, introducing self-contained execution contexts (SCECs) for reproducible admission and utility-based memory weighting.
@@
  • We conduct extensive evaluations on FEVER and HotpotQA benchmarks, demonstrating that SEDM consistently improves task accuracy while significantly reducing token overhead.
@@ 2 Related Work
Recent efforts in building self-evolving agents have focused on enabling systems to improve their reasoning or behavior over time without explicit retraining. Approaches such as Reflexion Shinn et al. (2023) and Voyager Wang et al. (2023) allow agents to iteratively refine their strategies by leveraging self-reflection and accumulated trajectories. Similarly, MEMIT Meng et al. (2022) demonstrates the feasibility of localized knowledge editing within large language models, suggesting a pathway for agents to evolve by continuously updating their internal representations. These studies highlight the importance of mechanisms that support autonomous adaptation and progressive self-improvement in dynamic environments.
@@
In parallel, research on agent memory has investigated how to store, retrieve, and utilize knowledge efficiently across long-horizon interactions. Episodic memory systems, such as those proposed by Park et al. Park et al. (2023), emulate human-like memory consolidation to support consistent long-term behavior in simulated social environments. Memory-augmented neural networks Graves et al. (2016) and differentiable neural dictionaries Kaiser et al. (2017) further demonstrate how structured memory access can enhance reasoning and generalization. More recently, retrieval-augmented generation frameworks tailored for interactive agents Mialon et al. (2023) have shown that dynamically grounding responses in curated external memories improves both interpretability and task success. Together, these works underscore the need for memory systems that are not only scalable but also adaptive to the agent’s evolving operational context.
@@ 3.3 Self-Scheduling in the Memory Controller
The memory controller manages and optimizes the repository through a self-scheduling policy. Unlike traditional systems that depend on costly per-query reranking Nogueira and Cho (2019); Ren et al. (2021); Ni et al. (2022), our approach establishes an evidence-based mechanism for both selecting memory items during retrieval and continuously refining the repository. It comprises two core functions: retrieval-time scheduling, which determines how to use memories effectively for an incoming query, and consolidation and progressive evolution, which curates a compact, high-quality memory set. Together, these components ensure that memory usage is grounded in verifiable utility signals, improving both efficiency and performance.
@@ 3.3.1 Retrieval-time Scheduling
The controller’s scheduling policy relies on a ranking signal aligned with realized utility, avoiding the instability and computational cost of on-the-fly large language model reranking Qin et al. (2023). Prior approaches typically use vector similarity or ad-hoc prompt-based scoring, but semantic similarity alone does not guarantee actual task benefit, and repeated reranking adds latency and variance Karpukhin et al. (2020); Izacard and Grave (2021) or ad-hoc prompt-based scoring Nogueira and Cho (2019); Qin et al. (2023).
@@ 3.3.2 Consolidation and Progressive Evolution
Progressive evolution is achieved by tracking usage and outcome signals and applying conservative updates to utility weights over time. Items that are rarely retrieved or consistently fail to provide positive utility are gradually decayed, reducing their influence in future selections. Conversely, items that repeatedly yield positive gains across related contexts are promoted. Promotion increases their weights and, when consistency is observed across multiple items, may trigger abstraction into higher-level insights. Such abstractions enable representative entries to replace families of consistent low-level experiences Goyal et al. (2019); Parisotto et al. (2019). If observed outcomes diverge significantly from admission-time evidence, items are demoted or queued for cleanup. All updates are logged with provenance to ensure the evolution process remains auditable Buneman et al. (2001); Lebo et al. (2013); Chirigati et al. (2016).
@@
The merged entry preserves essential content while aggregating evidence from the originals, and its weight $w(m_{\text{merged}})$ is reconciled to reflect combined support without double counting. Contributing items are archived or soft-deleted, preserving provenance if needed. By collapsing near-duplicates into single representatives, the repository reduces retrieval noise and strengthens utility signals Broder et al. (1997); Manku et al. (2007); Charikar (2002); Zhang et al. (2019).
Through these routines, the controller maintains a concise but reliable set of memories. Redundant entries are consolidated Manku et al. (2007), stable positives are promoted Kaelbling et al. (1996), abstractions generalize recurring insights Goyal et al. (2019), and harmful items are isolated or removed Toneva et al. (2019). As a result, the repository remains aligned with realized utility, ensuring that retrieval decisions exploit empirical evidence rather than ad-hoc reranking.
@@ 3.4 Cross-Domain Knowledge Diffusion
The abstraction process is rule-governed and minimal. Entities and domain-specific terms are replaced with typed placeholders, retaining actionable task–action structures while removing non-essential detail Goyal et al. (2019); Lake et al. (2015); Liang et al. (2017). The result is a compact snippet suitable for direct injection and controlled comparisons. To minimize orchestration cost, abstraction is generated alongside the SCEC-based A/B assessment within the distributed pipeline Chirigati et al. (2016).
@@ 4.1 Experimental Setup
Datasets.

### v3

@@ Abstract
In long-term multi-agent systems, the accumulation of trajectories and historical interactions makes efficient memory management a particularly challenging task, with significant implications for both performance and scalability. Existing memory management methods typically depend on vector retrieval and hierarchical storage, yet they are prone to noise accumulation, uncontrolled memory expansion, and limited generalization across domains. To address these challenges, we present SEDM (Self-Evolving Distributed Memory), a verifiable and adaptive framework that transforms memory from a passive repository into an active, self-optimizing component. SEDM integrates verifiable write admission based on reproducible replay, a self-scheduling memory controller that dynamically ranks and consolidates entries according to empirical utility, and cross-domain knowledge diffusion that abstracts reusable insights to support transfer across heterogeneous tasks. Evaluations on benchmark datasets demonstrate that SEDM improves reasoning accuracy while reducing token overhead compared with strong memory baselines, and further enables knowledge distilled from fact verification to enhance multi-hop reasoning. The results highlight SEDM as a scalable and sustainable memory management design for open-ended multi-agent collaboration. The code will be released upon acceptance of this paper.
@@ 1 Introduction
In recent years, the rapid development of large-scale multi-agent systems (MAS) Wooldridge2009MAS ; Lowe2017MADDPG ; Foerster2016Comm ; cite2 ; Busoniu2008Survey has expanded their application in diverse domains, including collaborative reasoning, decision-making, and autonomous planning cite1 . A fundamental challenge in open-ended, long-term tasks lies in enabling agents to effectively manage, interpret, and reutilize information accumulated through continuous interactions with both peers and their environment cite2 . In the absence of effective memory management design, the sheer scale of historical interactions can easily overwhelm computational resources and compromise decision quality cite3 .
@@
In open-ended and long-term multi-agent tasks, each agent relies on its past memories, the observed states of other agents, and the current environment to make decisions for subsequent actions or responses cite4 . During continuous interaction between agents and their environment, the MAS gradually accumulates extensive logs of interactions, invocation trajectories, and high-level policy memories openai2024gpt4 . Such overwhelming amounts of information directly impact the efficiency and cost of decision-making, often leading to higher monetary costs and longer contextual requirements for inference cite6 . Therefore, designing an efficient and sustainable memory mechanism has become a critical issue for modern long-term multi-agent systems.
@@
Current methods primarily adopt vector retrieval and hierarchical memory structures to manage storage and retrieval efficiently cite8 . Vector retrieval Johnson2019FAISS ; cite34 ; cite33 ; cite14 ; Guo2016DRMM leverages semantic similarity to identify relevant entries, while hierarchical organization arranges information in layered structures according to abstraction levels cite9 . These approaches have shown promise in improving retrieval accuracy and managing memory scalability cite10 . However, in complex collaborative multi-agent tasks, their effectiveness diminishes, as the underlying assumptions of stability and linear growth do not hold cite11 . This gap between theoretical promise and practical performance highlights several critical limitations that hinder their long-term applicability.
@@
One major challenge is the inevitable accumulation of noise, which severely degrades retrieval quality cite12 . As the memory size expands without constraint, the system faces exponentially increasing computational costs in both retrieval and context construction cite33 . This not only reduces overall efficiency but also amplifies the interference caused by redundant information cite14 . In particular, the presence of low-value or semantically irrelevant entries dilutes the contribution of high-quality information in retrieval results, impairing downstream task performance and leading to measurable declines in metrics cite15 . In addition, the cumulative noise effect increases response latency and accelerates the nonlinear consumption of computational and storage resources cite34 , ultimately threatening both scalability and stability in long-term MAS operations cite17 .
@@
To overcome these limitations, we introduce Scalable Self-Evolving Distributed Memory (SEDM), a framework that transforms memory from a passive repository into an adaptive, self-optimizing, and verifiable component for multi-agent systems. Unlike conventional designs that treat memory as a static store, SEDM continually refines knowledge to enhance learning and decision-making efficiency in dynamic task environments. It operationalizes memory as an active mechanism by integrating verifiability and continuous self-improvement into the memory lifecycle. At its core, memory items undergo a rigorous admission process based on self-contained execution contexts (SCECs), such as Docker and ReproZip Merkel2014Docker ; chirigati2016reprozip , which package all necessary information for environment-free replay and offline validation. This mechanism provides empirical evidence for utility at write time, ensuring that only useful, high-quality experiences enter the memory repository. Once admitted, memory items are dynamically managed by a self-scheduling controller and enhanced through cross-domain knowledge diffusion. The controller leverages admission-derived weights, combined with semantic similarity, to schedule retrieval-time usage without costly reranking, while consolidation and progressive evolution continuously refine the repository by promoting stable items, merging redundancies, and pruning harmful ones. Beyond single-task settings, SEDM abstracts reusable insights into general forms, enabling knowledge distilled in one domain to be safely transferred and re-validated in others. Together, these components establish a scalable and auditable memory mechanism that enhances reasoning accuracy, reduces overhead, and supports sustainable long-term multi-agent collaboration.
@@
We evaluate SEDM on two representative benchmarks, FEVER thorne-etal-2018-fever for fact verification and HotpotQA yang2018hotpotqadatasetdiverseexplainable for multi-hop reasoning, comparing against no-memory and G-Memory baselines zhang2025gmemorytracinghierarchicalmemory . The results show that SEDM consistently improves task accuracy while significantly reducing token overhead, thereby achieving a better balance between performance and efficiency. Ablation studies confirm that both the verifiable admission mechanism and the self-scheduling controller contribute progressively to this gain, with the latter playing a key role in constraining prompt growth without sacrificing accuracy. Furthermore, cross-domain evaluation demonstrates that memory distilled from one dataset can transfer to another, with factual knowledge from FEVER notably boosting performance on HotpotQA. These findings highlight SEDM as a scalable, adaptive, and generalizable memory framework for long-term multi-agent reasoning.
@@
  • We propose Self-Evolving Distributed Memory (SEDM), a novel framework that transforms memory from a passive repository into an adaptive, verifiable, and continuously improving component, introducing self-contained execution contexts (SCECs) for reproducible admission and utility-based memory weighting.
@@
  • We conduct extensive evaluations on LoCoMo, FEVER, and HotpotQA benchmarks, demonstrating that SEDM consistently improves task accuracy while significantly reducing token overhead.
@@ 2 Related Work
Recent efforts in building self-evolving agents have focused on enabling systems to improve their reasoning or behavior over time without explicit retraining. Approaches such as Reflexion shinn2023reflexion and Voyager wang2023voyager allow agents to iteratively refine their strategies by leveraging self-reflection and accumulated trajectories. Similarly, MEMIT meng2022memit demonstrates the feasibility of localized knowledge editing within large language models, suggesting a pathway for agents to evolve by continuously updating their internal representations. These studies highlight the importance of mechanisms that support autonomous adaptation and progressive self-improvement in dynamic environments.
@@
In parallel, research on agent memory has investigated how to store, retrieve, and utilize knowledge efficiently across long-horizon interactions. Episodic memory systems, such as those proposed by Park et al. park2023generativeagents , emulate human-like memory consolidation to support consistent long-term behavior in simulated social environments. Memory-augmented neural networks graves2016hybrid and differentiable neural dictionaries kaiser2017learning further demonstrate how structured memory access can enhance reasoning and generalization. More recently, retrieval-augmented generation frameworks tailored for interactive agents mialon2023augmented have shown that dynamically grounding responses in curated external memories improves both interpretability and task success. Together, these works underscore the need for memory systems that are not only scalable but also adaptive to the agent’s evolving operational context.
@@ 3.3 Self-Scheduling in the Memory Controller
The memory controller manages and optimizes the repository through a self-scheduling policy. Unlike traditional systems that depend on costly per-query reranking nogueira2019passage ; ren2021rocketqav2 ; ni2022large , our approach establishes an evidence-based mechanism for both selecting memory items during retrieval and continuously refining the repository. It comprises two core functions: retrieval-time scheduling, which determines how to use memories effectively for an incoming query, and consolidation and progressive evolution, which curates a compact, high-quality memory set. Together, these components ensure that memory usage is grounded in verifiable utility signals, improving both efficiency and performance.
@@ 3.3.1 Retrieval-time Scheduling
The controller’s scheduling policy relies on a ranking signal aligned with realized utility, avoiding the instability and computational cost of on-the-fly large language model reranking wu2023llmrerank . Prior approaches typically use vector similarity or ad-hoc prompt-based scoring, but semantic similarity alone does not guarantee actual task benefit, and repeated reranking adds latency and variance cite34 ; cite33 or ad-hoc prompt-based scoring nogueira2019passage ; wu2023llmrerank .
@@ 3.3.2 Consolidation and Progressive Evolution
Progressive evolution is achieved by tracking usage and outcome signals and applying conservative updates to utility weights over time. Items that are rarely retrieved or consistently fail to provide positive utility are gradually decayed, reducing their influence in future selections. Conversely, items that repeatedly yield positive gains across related contexts are promoted. Promotion increases their weights and, when consistency is observed across multiple items, may trigger abstraction into higher-level insights. Such abstractions enable representative entries to replace families of consistent low-level experiences goyal2019abstraction ; parisotto2019stabilizing . If observed outcomes diverge significantly from admission-time evidence, items are demoted or queued for cleanup. All updates are logged with provenance to ensure the evolution process remains auditable buneman2001provenance ; groth2012provenance ; chirigati2016reprozip .
@@
The merged entry preserves essential content while aggregating evidence from the originals, and its weight $w(m_{\text{merged}})$ is reconciled to reflect combined support without double counting. Contributing items are archived or soft-deleted, preserving provenance if needed. By collapsing near-duplicates into single representatives, the repository reduces retrieval noise and strengthens utility signals broder1997resemblance ; manku2007detecting ; charikar2002simhash ; zhang2019duplicate .
Through these routines, the controller maintains a concise but reliable set of memories. Redundant entries are consolidated manku2007detecting , stable positives are promoted kaelbling1996rlsurvey , abstractions generalize recurring insights goyal2019abstraction , and harmful items are isolated or removed toneva2019catastrophic . As a result, the repository remains aligned with realized utility, ensuring that retrieval decisions exploit empirical evidence rather than ad-hoc reranking.
@@ 3.4 Cross-Domain Knowledge Diffusion
The abstraction process is rule-governed and minimal. Entities and domain-specific terms are replaced with typed placeholders, retaining actionable task–action structures while removing non-essential detail goyal2019abstraction ; lake2015human ; dong2019neural . The result is a compact snippet suitable for direct injection and controlled comparisons. To minimize orchestration cost, abstraction is generated alongside the SCEC-based A/B assessment within the distributed pipeline chirigati2016reprozip .
@@ 4.1 Experimental Setup
4.1.1 Dataset and Model
Evaluation is conducted on the LoCoMo benchmark maharana2024LoCoMo , which consists of two components: (i) multi-turn dialogues (approximately 600 turns per dialogue, averaging 26,000 tokens) and (ii) question-answer (QA) pairs grounded in those dialogues. Each dialogue contains roughly 200 questions spanning single-hop rajpurkar2016squad , multi-hop yang2018hotpotqa , open-domain kwiatkowski2019natural , and temporal reasoning chen2021timeqa . All experiments are carried out with gpt-4o-mini to ensure consistency and comparability across evaluations.
Performance is measured using two complementary metrics: Token-level F1 Score (F1) rajpurkar2016squad , which captures the overlap between predicted and ground-truth answers, and BLEU-1 (B1) papineni2002bleu , which evaluates unigram-level lexical similarity.
4.1.2 Baselines
To evaluate the effectiveness of our framework SEDM, we compare it with several representative baselines for multi-session dialogue reasoning: LoCoMo maharana2024LoCoMo , a benchmark for assessing long-range retrieval and reasoning in multi-session conversations; ReadAgent lee2024human , a human-in-the-loop agent for long-context reading; MemoryBank zhong2024memorybank , a memory-augmented retrieval model; MemoryGPT packer2024memgptllmsoperatingsystems , an LLM operating system with hierarchical memory modules; A-Mem xu2025amemagenticmemoryllm , a dynamic agentic memory system that creates, links, and updates structured memories; Zep rasmussen2025zeptemporalknowledgegraph , a retrieval-based agent with structured memory access for temporally extended queries; LangMem wang2023augmentinglanguagemodelslongterm , an open-source framework connecting memory chains across sessions; Mem0 chhikara2025mem0buildingproductionreadyai , a modular memory system with explicit in-context memory operations; and G-memory zhang2025gmemorytracinghierarchicalmemory , which leverages hierarchical tracing for memory retrieval.
4.2 Main results
Table 1: Evaluation results on the LoCoMo benchmark dataset comparing SEDM with other memory-enabled systems. Models are evaluated on F1 rajpurkar2016squad and BLEU-1 (B1) papineni2002bleu across Single Hop rajpurkar2016squad , Multi-Hop yang2018hotpotqa , Open Domain kwiatkowski2019natural , Temporal, and Adversial questions chen2021timeqa . Higher is better. The top two best results are marked in Bold.
Method Single Hop Multi-Hop Open Domain Temporal Adversial
[…]
LoCoMo maharana2024LoCoMo 25.0 19.8 18.4 14.8 40.4 29.1 12.0 11.2 69.2 68.8
ReadAgent lee2024human 9.2 6.5 12.6 8.9 9.7 7.7 5.3 5.1 9.8 9.0
MemoryBank zhong2024memorybank 5.0 4.8 9.7 7.0 6.6 5.2 5.6 5.9 7.4 6.5
MemoryGPT packer2024memgptllmsoperatingsystems 26.7 17.7 25.5 19.4 41.0 34.3 9.2 7.4 43.2 42.7
A-Mem xu2025amemagenticmemoryllm 27.0 20.1 45.9 36.7 44.7 37.1 12.1 12.0 50.0 49.5
Zep rasmussen2025zeptemporalknowledgegraph 30.2 17.2 15.0 11.6 26.7 18.4 3.5 2.7 22.6 15.1
LangMem langchain2024 22.4 15.2 18.7 16.0 31.6 23.9 27.8 21.5 28.3 21.3
Mem0 chhikara2025mem0buildingproductionreadyai 27.3 18.6 18.6 13.9 34.0 24.8 26.9 21.1 30.4 22.2
G-memory zhang2025gmemorytracinghierarchicalmemory 34.6 26.6 9.05 7.2 53.5 44.0 32.4 25.6 11.3 9.3
SEDM (Ours) 33.5 24.4 12.1 9.2 51.7 37.0 47.5 33.1 12.1 9.3
Table 1 presents the results of SEDM compared with strong baselines, including LoCoMo, ReadAgent, MemoryBank, MemoryGPT, A-Mem, Zep, LangMem, Mem0, and G-memory, on the LoCoMo benchmark across Single Hop, Multi-Hop, Open Domain, Temporal, and Adversarial reasoning tasks.
Overall, the results show that SEDM delivers strong and consistent improvements in challenging reasoning scenarios. In particular, SEDM achieves the highest F1 and BLEU-1 scores on the Open Domain and Temporal settings, surpassing all other baselines by a substantial margin. For instance, SEDM improves Temporal reasoning by more than $15$ F1 points compared with the strongest baseline (G-memory), demonstrating its ability to capture and utilize temporally structured information effectively.
On the Single Hop setting, SEDM remains highly competitive, ranking among the top two models alongside G-memory, while showing significantly better balance across different question types. Although A-Mem and MemoryGPT achieve strong results on Multi-Hop and Adversarial tasks respectively, their performance drops markedly on other reasoning categories, whereas SEDM maintains stable and robust performance across the board.
These results highlight that the learned selective memory mechanism in SEDM provides clear advantages over heuristic, retrieval-based, and hierarchical memory systems. By selectively integrating context across sessions, SEDM enables LLMs to reason more effectively over long, multi-session dialogues and demonstrates strong generalization across diverse reasoning categories.
4.3 Efficient analysis on Fever and HotpotQA